What will be the pH value when we add 10 mL of 2 molar NaOH solution to 1 litre of 0.2 molar acetate buffer whose pH is 4.7? Pka = 4.7
Answer #1


pH of acidic buffer = pka + log(ch3cooNa/CH33COOH)

pka of CH3COOH = -logKa = 4.74

Total no of mol of buffer = 0.2*1 = 0.2 mole

no of mol of CH3COOH = x mole

No of mol of CH3COONa = NaOH = 0.2- x mole

4.74 = 4.74+log((0.2-x)/x)

   x = 0.1

no of mol of CH3COOH = x = 0.1 mole

No of mol of CH3COONa = NaOH = 0.2- 0.1 = 0.1 mole

No of mol of NaOH added = 10*2*10^-3 = 0.02 mole

pH = pka + log((CH3COONa+NaOH/CH3COOH-NaOH)

   = 4.74 + log((0.1+0.02)/(0.1-0.02))

pH = 4.92